# Spiking Neural Networks for Continuous Control via End-to-End Model-Based Learning

### v2

@@ 2.2 Leaky integrate-and-fire neuron model
Leaky integrate-and-fire models form a broad class of neuron models that vary in complexity, parameterization, and numerical methods (Gerstner et al.,, 2014). Choosing the LIF neuron model for robotics and machine learning tasks is based on a suitable trade-off between computational efficiency and biological plausibility because its simple, interpretable dynamics enable fast simulation and stable gradient-based learning, while still capturing key temporal integration properties necessary for control and sequence-based tasks. Variants of the LIF model differ in whether they include explicit synaptic currents, refractory periods, or adaptive thresholds; here, we adopt a two-variable (synaptic + membrane) formulation with subtractive reset and no refractory period, as this provides both computational efficiency and sufficiently rich temporal traces to support surrogate gradient learning.
@@ 2.3 Mandatory Training Considerations
Non-differentiable spike functions are approximated with smooth surrogate gradient methods during the backward pass. We evaluated several commonly used surrogate functions (sigmoid, Gaussian Spike Yin et al., (2021), and Super Spike Zenke and Ganguli, (2018)) and investigated the impact of steepness $\beta$ of the function on gradient magnitude and learning stability (see Appendix B).
@@
SNNs (similar to classic recurrent neural networks) are sensitive to weight initialization, as only certain parameter regimes yield stable behavior and effective learning. We employ a fluctuation-driven initialization (Rossbroich et al.,, 2022), which incorporates LIF time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ into the scaling of synaptic weights, improving upon common ANN-based initialization methods. The original scheme was developed for SNNs in (temporal) classification tasks, which allows for a priori assessment of input firing rates to identify an appropriate weight scale factor $\nu$. In our case, we estimate $\nu$ empirically as the training data is not known at time of network initialization and the network input is continuous instead of precomputed spike patterns. Our networks require careful tuning of parameters to achieve stable membrane potentials and gradients throughout the networks (see Appendix C).
@@
The prediction and policy networks are trained iteratively using batches from a replay buffer. We adopt unroll-based training (truncated backpropagation through time) with $w=10$ environment warmup steps and teacher forcing for the transition model. The policy model is updated via autoregressive predictions through the frozen transition model. See Appendix E, F, and G.
@@
The gradient magnitude in our networks is influenced by several factors. The LIF time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ (as well as $\tau_{\text{ada}}$ for ALIF neurons) control the length of eligibility traces any given input spike event leaves on the membrane voltage of downstream neurons. Further, the number of simulation steps $r$, the number of SNN sub-steps per simulation step, and the total number of spike events $S$ collectively affect the risk of exploding or vanishing gradients in partially recurrent networks. Finally, the choice of surrogate gradient function and the steepness parameter $\beta$ scale the gradients directly. Jointly searching the parameter space quickly becomes infeasible due to the large number of degrees of freedom in this optimization problem. We found that using an optimizer like adam Kingma and Ba, (2014), which can dynamically adapt learning rates per parameter, proved effective in stabilizing training. As the gradient magnitudes for prediction and policy networks vary, we evaluated a number of different learning rates for each model (see Appendix A).
@@ 2.4 Optional Training Components
Scheduled Learning Rates
Instead of using a constant learning rate, we explored an exponential decay schedule (see Appendix A).
Learnable Time Constants
LIF time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ (as well as $\tau_{\text{ada}}$ for ALIF neurons) can be learned for each neuron individually. This allows the network to tune its temporal processing dynamics. For a detailed overview, see Appendix H.
Regularization
In SNNs regularization may be critical to reduce the bursting and dying networks effect, and also to improve communication efficiency Hübotter et al., (2021). We evaluated multiple regularization techniques to stabilize training and improve generalization. These include $L_{2}$ weight decay (Appendix J), network activity constraints (Appendix K and I), as well as action penalties (Appendix L) and random action space exploration (Appendix M).
Reducing Network Parameters
Full rank connectivity between layers can quickly lead to a large number of parameters to optimize. We investigated a method to restrict the latent dimension of the encoded information, which can reduce the number of parameters needed in the networks by several orders of magnitude (see Appendix N).
@@ 2.5 Experimental setup
We evaluated Pred-Control SNNs on two simulated continuous control environments, a 2D planar arm and a 6-DOF robotic arm for a standard reaching task. To evaluate model performance in addition to looking at losses and gradients we also recorded the metrics listed in Table 2.
@@
A simple planar 2-joint robot arm is simulated using PyGame with the goal to reach for randomly generated target locations. The state includes joint angles and end-effector position, while the action specifies joint accelerations. The arm segment lengths are given as 0.5 and 0.4 arbitrary units (au) and a reaching attempt qualifies as successful when the end effector arrives within 0.05 au of the target. This task enables fast experimentation and analysis due to its short rollout time and low computational requirements. The unrestricted robot joint angles are given as both $\sin$ and $\cos$ of their angle, as well as the $x$ and $y$ coordinates of the end effector and the target position. Both the state and action representations are scaled such that they lie within [1,1][-1,1]. This task uses partial observability as velocity or acceleration information is not directly available and observations contain some small amount of noise.
3D Robotic Manipulandum
The second task features a simulated 6-DOF robot arm (Franka Emika Panda robot) operating in 3D space, where the goal is again to reach target positions. The threshold for a successful reaching attempt is defined as 0.123 m. This is simulated using Nvidias IsaacSim and IsaacLab Mittal et al., (2023). Here, the action corresponds to joint torques. The 3D task introduces more complex dynamics in the forward model with higher degrees of freedom and joint rotation limits. By default, we have disabled the influence of gravity so that the models only need to learn the kinematic forward model. When enabling gravity, the task is made more complex as the networks now also have to learn the inertia matrix. The state representation includes the rotation of each joint, scaled between [-1, 1], as well as the $x$, $y$, and $z$ coordinates of the end-effector and the target.
@@
2D and 3D environments run for 200 time steps per episode while data is collected in parallel using 64 simulated environments. At the start of each episode, a random robot configuration is sampled to define the target end-effector position, and a new random configuration is selected for the initial state. The initial velocities are always set to zero.
For evaluation, the 2D task uses 8 hand-crafted initial and target configurations shared across runs. In the 3D setting, we sample 64 fixed random initial and target states per run. Although these vary with the seed and prevent exact comparisons between runs, they offer internal consistency.
Iterative Development Process
[…]
Table 2: Overview of task performance metrics. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are preferred.
@@
Time to Target Time step at which the end-effector first reaches the target zone. $\downarrow$
@@
Unrolled State MSE Mean squared error of the prediction model during autoregressive rollout. $\downarrow$
Mean Active Neurons Average number of neurons that spike at least once per episode, averaged across all layers. $\uparrow$
@@ 3 Results
Figure 3: Impact of optional model components on task performance. Shown are four models solving the 6-DOF robotic arm reaching task. The first model uses hyperparameters from the 2D reaching task (dark blue) without optional model components. The Pred-Control SNN has re-tuned initialization parameters ($\tau_{\text{mem}}$, $\tau_{\text{syn}}$, and $\nu$) and undergoes ablation of optional components. Results show a clear improvement after retuning and after including learnable time constants, ALIF neurons, and latent compression. The full Pred-Control SNN outperforms all ablated models despite also needing to learn the effects of gravity on dynamics.
As shown in Figure 3, hyperparameters tuned for the 2D planar reaching task fail to generalize to the 3D setting. Even applying hyperparameters from an intermediate 4-DOF setup does not yield satisfactory performance. This suggests that SNNs for control are sensitive to task dynamics and require careful re-tuning of key parameters—most notably the synaptic and membrane time constants ($\tau_{\text{syn}}$, $\tau_{\text{mem}}$) and the weight scaling factor $\nu$. These govern both the temporal response properties of the network and the stability of gradient propagation during training.
@@
We find that adding learnable time constants consistently improves performance, especially when paired with appropriately scaled learning rates (see Appendix H). This adaptation enables the model to fine-tune its intrinsic temporal dynamics in response to task demands, partly mitigating the sensitivity to initialization. Notably, we found that assigning higher learning rates to time constants than to weights and biases improves convergence speed and stability without destabilizing training.
@@
Replacing the LIF neurons with adaptive LIF (ALIF) units further enhances performance (see Appendix I). Our ALIF neurons incorporate two key mechanisms: I) a linearly decaying threshold to force spiking behavior and II) an adaptive component that increases with recent activity and gradually relaxes back to baseline. This mechanism provides dual benefits: it enables previously inactive neurons to re-engage in encoding and learning, and it prevents persistent overactivation of units during long episodes. Together, these dynamics facilitate better credit assignment over time and help stabilize network activity during training and inference.
We also observe performance gains from introducing latent-space compression between the spiking layers (see Appendix N). This method reduces the dimensionality of the latent representation passed between network stages, freeing up capacity to increase the number of spiking neurons per layer while maintaining a fixed parameter budget. In our best-performing models, compressed architectures used 2048 spiking neurons per layer and a 64-dimensional latent bottleneck, compared to 512-neuron layers in the uncompressed case. Both models have roughly the same number of learned parameters overall. We hypothesize that the higher neuron count improves precision in encoding and decoding continuous-valued signals, while the compression enforces a compact, task-relevant representation that aids generalization.
@@
Figure 4: Pred-Control SNN behavior during task execution. Shown are 3D trajectories and Euclidean error traces (I) over the course of training, along with spiking activity (II) and voltage traces (III) from the prediction and policy networks. The model exhibits progressively smoother control and more consistent activity as training proceeds.
Figure 4 illustrates the evolution of network behavior during training. The 3D trajectory of the robotic arm becomes more accurate and stable over time, with distance-to-target errors decreasing steadily across epochs. Simultaneously, the membrane voltages and spike trains of both the prediction and policy networks stabilize, indicating that the network develops a robust internal representation of both control dynamics and task structure. We interpret this as a sign that the model transitions from exploration to a settled control regime.
These experiments confirm that fully spiking networks, when trained with deep learning–inspired methods adapted to the spiking domain, can solve continuous, high-dimensional motor tasks with high accuracy and stability. Altogether, our findings demonstrate that SNNs can match the functional demands of complex control settings, opening the door to scalable, adaptable, and energy-efficient spiking systems for robotics and beyond.
@@ 4 Discussion and Conclusion
Spiking neural networks offer a compelling model of temporal computation for embodied systems, drawing inspiration from biological principles while promising advantages in energy efficiency and dynamic signal processing. Previous work has demonstrated that SNNs can be applied to continuous control, often via equation-based controllers (Slijkhuis et al.,, 2023; Agliati et al.,, 2025) or local adaptation frameworks such as the Neural Engineering Framework (DeWolf et al.,, 2016; Iacob et al.,, 2020; Marrero et al.,, 2024). Yet these approaches typically remain limited to low-dimensional tasks or constrained by analytical assumptions. Our work demonstrates that, when equipped with modern training tools adapted from deep learning, spiking architectures like the Pred-Control SNN can scale to high-dimensional robotic control. To our knowledge, this is the first demonstration of a fully spiking, end-to-end trainable predictive-control architecture operating in high-DoF continuous robotic arm tasks, helping to close a long-standing gap between biologically inspired modeling and practical machine learning. This positions our study beyond earlier demonstrations of spiking controllers, showing that deep-learning–style methods enable SNNs to function as competitive continuous controllers rather than as proof-of-concept demonstrations.
@@
By training our Pred-Control SNN using surrogate gradients and evaluating a suite of architectural extensions, we systematically identified several components that significantly affect performance and learning dynamics. Adaptive thresholds improved neuron participation and supported sparse activity patterns when well-tuned, consistent with prior findings on adaptive spiking neurons in classification settings (Yin et al.,, 2021). Decaying thresholds enabled silent neurons to reactivate over time, serving as an effective alternative to static lower-bound activity constraints. Spike regularization reduced bursting but required careful calibration to avoid performance loss. Learning time constants, especially on a per-neuron basis in log space, offered flexibility to compensate for poor initialization, echoing stabilization strategies such as fluctuation-driven initialization (Rossbroich et al.,, 2022), provided that their learning rates were sufficiently high. Additionally, loss shaping and latent space compression translated well from ANN training, aiding credit assignment and signal precision within the constraints of sparse spiking dynamics.
@@
Together, our results confirm that many principles of deep learning, including dynamic adaptation, sparsity control, and structured dimensionality reduction, are not only compatible with spiking architectures but may be essential for unleashing their potential in continuous tasks. They offer a reproducible blueprint for making SNNs robust, trainable, and effective in real-valued motor control. This moves beyond earlier demonstrations in arm reaching with STDP or NEF rules (Fernández et al.,, 2021; Juarez-Lora et al.,, 2022; DeWolf et al.,, 2016), neuromorphic hardware pilots (Zhao et al.,, 2020; DeWolf et al.,, 2023; Paredes-Vallés et al.,, 2024), or policy-focused surrogate gradient RL (Tang et al.,, 2021; Park et al.,, 2025; Oikonomou et al.,, 2023; Zanatta et al.,, 2024). In contrast, our predictive-control architecture explicitly integrates forward modeling with control, positioning this work as a step toward scalable spiking systems that unify deep learning training with predictive computation. For the *Neural Computation* audience, this suggests that SNNs can move from classification benchmarks toward the kind of adaptive, embodied computation that biological systems perform.
@@
Our study has limitations worth noting. First, the reliance on backpropagation-through-time introduces significant computational overhead in both memory and runtime, an issue exacerbated in SNNs due to their internal sub-timesteps and fine-grained temporal resolution (Neftci et al.,, 2019; Zenke and Vogels,, 2021). This imposes a bottleneck for scaling to longer task horizons, real-time execution, or complex multi-agent scenarios, and represents a central obstacle to broader applicability. Second, while our models generalize across varying initial conditions, they remain sensitive to hyperparameters such as regularization strength and adaptation rates. From a robotics or ML perspective this is a limitation, but from a neural computation perspective it may be better viewed as a result: biological systems also rely on finely tuned dynamics to maintain robustness. Future work should explore additional homeostatic mechanisms inspired by biological regulation, akin to threshold adaptation or learnable time constants. Likewise, incorporating structured connectivity patterns such as partially inhibitory lateral connections or sparsity-promoting weight matrices may enable decorrelation across neurons while preserving the benefits of distributed population codes. Finally, most existing SNN control work, including our own, currently uses rate coding for motor output and state representation, leaving the potential of alternative spike coding schemes for control largely untapped (Slijkhuis et al.,, 2023).

### v3

@@ 2.2 Leaky integrate-and-fire neuron model
Leaky integrate-and-fire models form a broad class of neuron models that vary in complexity, parameterization, and numerical methods (Gerstner et al., 2014). Choosing the LIF neuron model for robotics and machine learning tasks is based on a suitable trade-off between computational efficiency and biological plausibility because its simple, interpretable dynamics enable fast simulation and stable gradient-based learning, while still capturing key temporal integration properties necessary for control and sequence-based tasks. Variants of the LIF model differ in whether they include explicit synaptic currents, refractory periods, or adaptive thresholds; here, we adopt a current-based leaky integrate-and-fire (LIF) neuron with explicit synaptic current dynamics with subtractive reset and no refractory period, as this provides both computational efficiency and sufficiently rich temporal traces to support surrogate gradient learning. While a single-variable LIF model may suffice for this task, we adopt a current-based LIF with synaptic dynamics to remain compatible with the fluctuation-driven initialization of Rossbroich et al. (2022), avoiding re-derivation for a different neuron model.
@@
The complete LIF model dynamics are summarized in Algorithm 1 and visualized in Figure 2 for different settings of time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$.
@@ 2.3 Mandatory Training Considerations
Non-differentiable spike functions are approximated with smooth surrogate gradient methods during the backward pass. We evaluated several commonly used surrogate functions (sigmoid, Gaussian Spike Yin et al. (2021), and Super Spike Zenke and Ganguli (2018)) and investigated the impact of steepness $\beta$ of the function on gradient magnitude and learning stability (see Appendix B).
@@
SNNs (similar to classic recurrent neural networks) are sensitive to weight initialization, as only certain parameter regimes yield stable behavior and effective learning. We employ a fluctuation-driven initialization (Rossbroich et al., 2022), which incorporates LIF time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ into the scaling of synaptic weights, improving upon common ANN-based initialization methods. The original scheme was developed for SNNs in (temporal) classification tasks, which allows for a priori assessment of input firing rates to identify an appropriate weight scale factor $\nu$. In our case, we estimate $\nu$ empirically as the training data is not known at time of network initialization and the network input is continuous instead of precomputed spike patterns. Our networks require careful tuning of parameters to achieve stable membrane potentials and gradients throughout the networks (see Appendix C).
@@
The prediction and policy networks are trained iteratively using batches from a replay buffer. We adopt unroll-based training (truncated backpropagation through time) with $w=10$ environment warmup steps and teacher forcing for the transition model. The policy model is updated via autoregressive predictions through the frozen transition model Unless stated otherwise, networks are trained using mini-batches of 256 samples for 25 batches per iteration. See Appendix E, F, and G.
@@
The gradient magnitude in our networks is influenced by several factors. The LIF time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ (as well as $\tau_{\text{ada}}$ for ALIF neurons) control the length of eligibility traces any given input spike event leaves on the membrane voltage of downstream neurons. Further, the number of simulation steps $r$, the number of SNN sub-steps per simulation step, and the total number of spike events $S$ collectively affect the risk of exploding or vanishing gradients in partially recurrent networks. Finally, the choice of surrogate gradient function and the steepness parameter $\beta$ scale the gradients directly. Jointly searching the parameter space quickly becomes infeasible due to the large number of degrees of freedom in this optimization problem. We found that using the adam optimizer Kingma and Ba (2014), which can dynamically adapt learning rates per parameter, proved effective in stabilizing training. As the gradient magnitudes for prediction and policy networks vary, we evaluated a number of different learning rates for each model (see Appendix A). The final learning rates used are $\alpha_{\tau}=0.01$ for time constants and $\alpha=0.001$ for all other parameters across both networks without decay.
@@ 2.4 Optional Training Components
Several optional architectural and training components were evaluated to improve stability, temporal credit assignment, and parameter efficiency in spiking predictive control. Learnable membrane, synaptic, and adaptation time constants allow neurons to tune their intrinsic temporal dynamics and substantially improve convergence and final performance (Appendix H). Adaptive LIF (ALIF) neurons introduce activity-dependent thresholds that stabilize firing rates, re-engage silent neurons, and improve long-horizon credit assignment (Appendix I). Latent-space compression constrains the dimensionality of inter-layer representations, enabling wider spiking layers at fixed parameter budgets and yielding favorable performance–efficiency trade-offs (Appendix N). Additional components, including scheduled learning rates (Appendix A), weight decay (Appendix J), activity regularization (Appendix K), action penalties (Appendix L), and exploratory action noise (Appendix M), were explored but did not consistently improve final performance beyond the components above and are therefore omitted from the final model.
@@ 2.5 Experimental setup
We evaluated Pred-Control SNNs on two simulated continuous control environments, a 2D planar arm and a 6-DOF robotic arm. Both tasks are goal-reaching problems trained end-to-end using identical data collection, unrolling, and optimization pipelines. Performance is assessed using cumulative distance to target, success rate, time on target, and prediction mean squared error (Prediction MSE). Prediction MSE is computed from autoregressive rollouts of the learned forward model and reflects long-horizon prediction accuracy rather than one-step prediction loss (see Table 2).
@@
The 2D task simulates a two-joint planar arm with fixed segment lengths in PyGame. The state includes joint angles and end-effector position, while actions specify joint accelerations. Observations are partially observable (no velocity information in state) and include mild noise, enabling fast experimentation and hyperparameter screening. The unrestricted robot joint angles are given as both $\sin$ and $\cos$ of their angle, as well as the $x$ and $y$ coordinates of the end effector and the target position. Both the state and action representations are scaled such that they lie within [1,1][-1,1].
3D Robot Arm (6-DOF)
The 3D task uses a simulated Franka Emika Panda robot with six actuated degrees of freedom and torque control, implemented in NVIDIA Isaac Sim / Isaac Lab Mittal et al. (2023). The goal is to reach target positions in Cartesian space within a fixed distance threshold. All 3D experiments reported in this work are conducted with gravity enabled. The state representation includes joint rotations (scaled to [1,1][-1,1]), end-effector position, and target position. The threshold for a successful reaching attempt is defined as 0.123 m. The 3D task introduces more complex dynamics in the forward model with higher degrees of freedom and joint rotation limits.
@@
We fix the simulation time step at $\Delta t=0.02$ s (50 Hz) for both the 2D and 3D tasks. 2D and 3D episodes last 200 environment steps while data is collected in parallel across 64 simulated environments. At the start of each episode, a random robot configuration is sampled to define the target end-effector position, and a new random configuration is selected for the initial state. The initial velocities are always set to zero.
For evaluation, the 2D task uses 8 hand-crafted initial and target configurations shared across runs. In the 3D setting, we sample 64 fixed random initial and target states per run. Although these vary with the seed and prevent exact comparisons between runs, they offer internal consistency. Unless stated otherwise, x-axes in learning curves refer to training iterations rather than environment or network time steps.
Hyperparameter search
To identify a robust training configuration, we first applied the necessary methods from Section 2.3 to the 2D reaching task, to identify viable time constants ($\tau_{\text{mem}}$, $\tau_{\text{syn}}$, and $\tau_{\text{ada}}$ for ALIF neurons), weight scalings (input, output, recurrent proportion $\rho$, and expected pre-synaptic spike rate $\nu$), learning rates ($\alpha_{\bm{\upsilon}}$ and $\alpha_{\bm{\pi}}$), surrogate gradient function (type and scale $\beta$), network architecture, batch size ($n$), memory capacity ($M$), and unroll steps ($T_{\bm{\upsilon}}$ and $T_{\bm{\pi}}$), as well as the optional methods outlined in Section 2.4. Hyperparameters optimized for the 2D task did not transfer reliably to the 3D setting, requiring task-specific retuning of temporal parameters and weight scalings. The search results are shown in the Appendix and all final hyperparameters used in the reported experiments are summarized in Appendix R.
Table 2: Overview of task performance metrics (top) and network spiking statistics (bottom). Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are preferred.
@@
Prediction MSE Mean squared error of the prediction model during autoregressive rollout. $\downarrow$
Mean Active Neurons Number of neurons that spike at least once per episode, averaged across all layers. $\uparrow$
@@ 3 Results
We compare four model classes in the main results: (i) a full Pred-Control SNN (large variant, 2048 neurons per layer, $\approx$0.9M parameters) combining learnable time constants, ALIF neurons, and latent-space compression; (ii) a basic SNN that disables these components (512 neurons per layer, $\approx$0.8M parameters); and non-spiking recurrent baselines with identical training pipelines but standard floating-point RNN modules, evaluated in both a small (S) (512 neurons per layer, $\approx$0.8M parameters) and large (L) (2048 neurons per layer, $\approx$12.7M parameters) configuration. All models are trained under identical data collection, unrolling, and optimization protocols. A more complete overview on baseline model performance is given in Appendix Q.
Figure 3: Main results on the 3D reaching task. Learning curves comparing a non-spiking recurrent baseline (small and large variants), a basic spiking controller, and the full Pred-Control SNN (large). Top row: task-level performance metrics (cumulative distance to target, success rate, time on target, and Prediction MSE). The Pred-Control SNN achieves task performance on par with the large non-spiking baseline despite using more than an order of magnitude fewer parameters, while clearly outperforming the basic SNN and the small non-spiking baseline. The non-spiking baselines converge slightly faster and achieve lower long-horizon prediction error, but this advantage does not translate into superior control performance. The Pred-Control SNN learns more gradually but exhibits smoother and more stable optimization dynamics. Curves show mean $\pm$ s.e.m. over 10 random seeds.
As shown in Figure 3, the full Pred-Control SNN reaches comparable cumulative distance, success rate, and time-on-target performance to the large non-spiking baseline, while using more than an order of magnitude fewer parameters. Both models substantially outperform the basic SNN, indicating that architectural extensions beyond a minimal spiking controller are required to solve the 3D task reliably. The small non-spiking baseline achieves reasonable control performance but remains consistently below both large models, highlighting the importance of model capacity for this task.
The non-spiking baselines achieve lower Prediction MSE during long-horizon autoregressive rollouts, reflecting more accurate forward predictions. However, this advantage does not translate into improved control performance. Despite higher Prediction MSE, the full Pred-Control SNN matches or exceeds the task-level metrics of the baselines. This suggests that effective control does not require a perfectly accurate predictive model, but rather a model that is sufficiently expressive to capture task-relevant dynamics and provide stable gradients for policy optimization.
@@
We find that adding learnable time constants consistently improves performance, especially when paired with appropriately scaled learning rates (see Appendix H and O). This adaptation enables the model to fine-tune its intrinsic temporal dynamics in response to task demands and improves both convergence and final performance, partly mitigating the sensitivity to initialization. Notably, we found that assigning higher learning rates to time constants than to weights and biases improves convergence speed and stability without destabilizing training.
@@
Replacing the standard LIF neurons with adaptive LIF (ALIF) units further enhances performance (see Appendix I and O). Our ALIF neurons incorporate two key mechanisms: I) a linearly decaying threshold to force spiking behavior and II) an adaptive component that increases with recent activity and gradually relaxes back to baseline. This mechanism provides dual benefits: it enables previously inactive neurons to re-engage in encoding and learning, and it prevents persistent overactivation of units during long episodes. Together, these dynamics facilitate better credit assignment over time and help stabilize network activity during training and inference.
[…]
@@
Figure 4: Pred-Control SNN behavior during task execution. Shown are 3D trajectories and Euclidean error traces (I) over the course of training, along with spiking activity (II) and voltage traces (III) from the prediction and policy networks. Top-row plots are shown over 200 environment time steps per episode. The middle and bottom rows are plotted over internal spiking model time, where each environment step corresponds to 7 spiking substeps, resulting in 1400 model time steps per episode. This distinction reflects the multi-timescale structure of the controller and is consistent across all experiments. The model exhibits progressively smoother control and more consistent activity as training proceeds. Spiking activity in the policy network becomes denser and more regular, while the prediction network shows and opposing trend. Mild oscillatory behavior near the target is visible in later training stages.
Figure 4 illustrates representative trajectories and internal network dynamics during task execution. As training progresses, end-effector trajectories become smoother and more consistent, and distance-to-target errors decrease steadily. Simultaneously, the membrane voltages and spike trains of both the prediction and policy networks stabilize, indicating that the network develops a robust internal representation of both control dynamics and task structure. Notably, as training stabilizes, spiking activity in the policy network becomes denser and more regular than in the prediction network, consistent with the policy’s role in producing continuous control signals rather than sparse state-change representations. We interpret this as a sign that the model transitions from exploration to a settled control regime. At later training stages, small oscillations around the target can be observed in the distance traces, despite high time-on-target performance. We further evaluate robustness to within-episode target shifts and observe rapid re-acquisition of perturbed targets without online learning (see Appendix P).
These experiments confirm that fully spiking networks, when equipped with adaptive temporal dynamics and efficient representations, can solve continuous, high-dimensional motor tasks with high accuracy and stability comparable to non-spiking recurrent baselines. Importantly, this is achieved with substantially fewer parameters and without requiring highly accurate long-horizon predictions, highlighting a key distinction between prediction fidelity and control effectiveness in model-based spiking systems. Altogether, our findings demonstrate that SNNs can match the functional demands of complex control settings, opening the door to scalable, adaptable, and energy-efficient spiking systems for robotics and beyond.
@@ 4 Discussion and Conclusion
Spiking neural networks offer a compelling model of temporal computation for embodied systems, drawing inspiration from biological principles while promising advantages in energy efficiency and dynamic signal processing. Previous work has demonstrated that SNNs can be applied to continuous control, often via equation-based controllers (Slijkhuis et al., 2023; Agliati et al., 2025) or local adaptation frameworks such as the Neural Engineering Framework (DeWolf et al., 2016; Iacob et al., 2020; Marrero et al., 2024). Yet these approaches typically remain limited to low-dimensional tasks or constrained by analytical assumptions. Our work demonstrates that, when equipped with modern training tools adapted from deep learning, spiking architectures like the Pred-Control SNN can scale to high-dimensional robotic control. To our knowledge, this is the first demonstration of a fully spiking, end-to-end trainable predictive-control architecture operating in high-DoF continuous robotic arm tasks, helping to close a long-standing gap between biologically inspired modeling and practical machine learning. This positions our study beyond earlier demonstrations of spiking controllers, showing that deep-learning–style methods enable SNNs to function as competitive continuous controllers rather than as proof-of-concept demonstrations.
@@
By training our Pred-Control SNN using surrogate gradients and evaluating a suite of architectural extensions, we systematically identified several components that significantly affect performance and learning dynamics. Adaptive thresholds improved neuron participation and supported sparse activity patterns when well-tuned, consistent with prior findings on adaptive spiking neurons in classification settings (Yin et al., 2021). Decaying thresholds enabled silent neurons to reactivate over time, serving as an effective alternative to static lower-bound activity constraints. Spike regularization reduced bursting but required careful calibration to avoid performance loss. Learning time constants, especially on a per-neuron basis in log space, offered flexibility to compensate for poor initialization, echoing stabilization strategies such as fluctuation-driven initialization (Rossbroich et al., 2022), provided that their learning rates were sufficiently high. Additionally, loss shaping and latent space compression translated well from ANN training, aiding credit assignment and signal precision within the constraints of sparse spiking dynamics. Viewed together, learnable time constants, adaptive thresholds, and latent-space compression address complementary aspects of temporal credit assignment in spiking predictive control. However, their combined effect enable stable, fast, and competitive spiking control.
A central insight emerging from our results is that prediction accuracy does not directly translate to control quality. While the non-spiking baselines achieve substantially lower long-horizon Prediction MSE, this advantage does not yield superior task-level performance. Instead, the Pred-Control SNN achieves comparable or better control despite less accurate autoregressive predictions. This suggests that effective control requires a predictive model that is sufficiently expressive to capture task-relevant dynamics and provide stable gradients for policy optimization, rather than one that minimizes prediction error in isolation. From a model-based control perspective, this aligns with the notion that approximate forward models can be entirely sufficient for planning and control when errors remain structured and bounded.
@@
Together, our results confirm that many principles of deep learning, including dynamic adaptation, sparsity control, and structured dimensionality reduction, are not only compatible with spiking architectures but may be essential for unleashing their potential in continuous tasks. They offer a reproducible blueprint for making SNNs robust, trainable, and effective in real-valued motor control. This moves beyond earlier demonstrations in arm reaching with STDP or NEF rules (Fernández et al., 2021; Juarez-Lora et al., 2022; DeWolf et al., 2016), neuromorphic hardware pilots (Zhao et al., 2020; DeWolf et al., 2023; Paredes-Vallés et al., 2024), or policy-focused surrogate gradient RL (Tang et al., 2021; Park et al., 2025; Oikonomou et al., 2023; Zanatta et al., 2024). In contrast, our predictive-control architecture explicitly integrates forward modeling with control, positioning this work as a step toward scalable spiking systems that unify deep learning training with predictive computation. For the neural computation audience, this suggests that SNNs can move from classification benchmarks toward the kind of adaptive, embodied computation that biological systems perform.
@@
Our study has limitations worth noting. First, the reliance on backpropagation-through-time introduces significant computational overhead in both memory and runtime, an issue exacerbated in SNNs due to their internal sub-timesteps and fine-grained temporal resolution (Neftci et al., 2019; Zenke and Vogels, 2021). This imposes a bottleneck for scaling to longer task horizons, real-time execution, or complex multi-agent scenarios, and represents a central obstacle to broader applicability. Second, while our models generalize across varying initial conditions, they remain sensitive to hyperparameters such as regularization strength and adaptation rates. From a robotics or ML perspective this is a limitation, but from a neural computation perspective it may be better viewed as a result: biological systems also rely on finely tuned dynamics to maintain robustness. Future work should explore additional homeostatic mechanisms inspired by biological regulation, akin to threshold adaptation or learnable time constants. Likewise, incorporating structured connectivity patterns such as partially inhibitory lateral connections or sparsity-promoting weight matrices may enable decorrelation across neurons while preserving the benefits of distributed population codes. Finally, most existing SNN control work, including our own, currently uses rate coding for motor output and state representation, leaving the potential of alternative spike coding schemes for control largely untapped (Slijkhuis et al., 2023).
Additionally, we observe mild oscillatory behavior near the target in some trajectories at later training stages for the Pred-Control SNN as well as the non-spiking baseline models. These oscillations occur despite high success rates and time-on-target performance and likely reflect a combination of local attractor dynamics exhibited by the policy network, limited prediction horizon, and the absence of strong action-smoothing penalties. Notably, the controller is not explicitly incentivized to come to rest at the target and the $\ell_{2}$ loss is weak close to the target position. We explored alternative loss formulations, including mixed $\ell_{1}/\ell_{2}$ distance penalties designed to increase attraction toward the target center, but these did not reliably eliminate oscillations (data not shown). Similarly, action regularization reduced overall movement speed and time-on-target but did not fully suppress oscillatory behavior unless set to levels that prevented motion altogether (see Appendix L), suggesting a trade-off between control smoothness and learning efficiency. This suggests that oscillations arise not from instability in the spiking dynamics, but from an underconstrained stopping criterion in continuous control via supervised learning, a challenge shared with non-spiking controllers. Future work could explore structured action regularizers, multi-rate control schemes, or predictive uncertainty estimates to mitigate such effects without compromising performance. An alternative approach to reaching tasks in the domain of reinforcement learning is known by applying more complex reward shaping compared to a simple distance based measure, where extra rewards might be granted close to the target, to give a stronger incentive (steeper loss landscape / stronger attractor) towards the target position.
Although spiking neural networks are often motivated by energy efficiency, our work does not explicitly optimize for sparsity or minimal spike counts. Under a rate-coding assumption, reducing spike activity necessarily trades representational precision for efficiency, which may or may not be desirable depending on task demands. For continuous motor control, population size and spike density jointly determine how finely continuous variables can be represented. In principle, equivalent control signals could be produced by larger populations with proportionally lower firing rates, or by discretized action spaces with fewer output levels. However, in the absence of explicit hardware constraints, sparsity becomes an ambiguous objective. We therefore adopt a hardware-ignorant perspective: rather than enforcing biological or neuromorphic limits, we focus on identifying architectural principles that enable robust learning and control. The reported spike activity metrics are thus interpreted as descriptive signals of network dynamics rather than direct proxies for energy consumption. Explicit energy optimization and deployment on neuromorphic hardware remain important directions for future work.